Kubin Company’s relevant range of production is 23,000 to 27,500 units. When it produces and sells 25,250 units, its average costs per unit are as follows:

  

Average Cost per Unit
Direct materials $ 8.30
Direct labor $ 5.30
Variable manufacturing overhead $ 2.80
Fixed manufacturing overhead $ 6.30
Fixed selling expense $ 4.80
Fixed administrative expense $ 3.80
Sales commissions $ 2.30
Variable administrative expense $ 1.80

Required:

1. For financial accounting purposes, what is the total amount of product costs incurred to make 25,250 units?

2. For financial accounting purposes, what is the total amount of period costs incurred to sell 25,250 units?

3. For financial accounting purposes, what is the total amount of product costs incurred to make 27,500 units?

4. For financial accounting purposes, what is the total amount of period costs incurred to sell 23,000 units?

(For all requirements, do not round intermediate calculations.)

1. Total amount of product costs
2. Total amount of period costs incurred
3. Total amount of product costs
4. Total amount of period costs
Answer #1
Part-1
Variable Product Costs per unit = Direct Materials + Direct Labor + Variable Manufacturing Overhead
Variable Product Costs per unit = $8.30 + $5.30 + $2.80
Variable Product Costs per unit = $16.40
Fixed Product Cost = Fixed Manufacturing Overhead per unit * Number of units
Fixed Product Cost = $6.30 * 25,250
Fixed Product Cost = $159075
Total product Cost= Variable Product Cost + Fixed Product Cost
$16.40X 25250+ 159075=$573175
Part-2
Variable Period Costs per unit = Sales Commissions + Variable Administrative Expense
Variable Period Costs per unit = $2.30 + $1.80
Variable Period Costs per unit = $4.10
Fixed Period Costs per unit = Fixed Selling Expense + Fixed Administrative Expense
Fixed Period Costs per unit = $4.80 + $3.80
Fixed Period Costs per unit = $8.60
Fixed Period Costs = Fixed Period Costs per unit * Number of units
Fixed Period Costs = $8.60 * 25250=$217150
Total Period Cost = Variable Period Cost+ Fixed period Cost
=($4.10X25250)+217150=$320675
Part-3
Variable Product Costs per unit = $16.40
Fixed Product Cost = $159075
Total Product Cost = $16.40 * 27500 + $159075
Total Product Cost = $610075
Part-4
Variable Period Costs per unit = $4.10
Fixed Period Costs = $217150
Total Period = $4.10 * 23,000 + $217150
Total Period = $311450